# Can We Verify Step by Step for Incorrect Answer Detection?

### v2

@@ Abstract
Chain-of-Thought (CoT) prompting has marked a significant advancement in enhancing the reasoning capabilities of large language models (LLMs). Previous studies have developed various extensions of CoT, which focus primarily on enhancing end-task performance. In addition, there has been research on assessing the quality of reasoning chains in CoT. This raises an intriguing question: Is it possible to predict the accuracy of LLM outputs by scrutinizing the reasoning chains they generate? To answer this research question, we introduce a benchmark, R2PE, designed specifically to explore the relationship between reasoning chains and performance in various reasoning tasks spanning five different domains. This benchmark aims to measure the falsehood of the final output of LLMs based on the reasoning steps. To make full use of information in multiple reasoning chains, we propose the process discernibility score (PDS) framework that beats the answer-checking baseline by a large margin. Concretely, this resulted in an average of $5.1\%$ increase in the F1 score and $2.97\%$ improvement in AUC-PR across all 45 subsets within R2PE. We further demonstrate our PDS’s efficacy in advancing open-domain QA accuracy. Data and code are available at https://github.com/XinXU-USTC/R2PE.
Can We Verify Step by Step for Incorrect Answer Detection?
Xin Xu, Shizhe Diao, Can Yang, Yang Wang Hong Kong University of Science and Technology xxuca@connect.ust.hk {sdiaoaa, macyang, yangwang}@ust.hk
Figure 1: An example from HotpotQA that GPT-4 outputs highly consistent but wrong answer. In this particular example, our PDS can detect conflicting information about the actor (colored by red) and make a prediction that aligns with the label, while the ADS that uses the answer agreement does not.
footnotetext: †Corresponding author.
@@ 1 Introduction
Recent development in large language models (LLMs) (Brown et al., 2020; OpenAI, 2023; Bubeck et al., 2023) has showcased their remarkable aptitude for tackling diverse downstream tasks. Given several demonstrations with reasoning steps, LLMs exhibit a formidable capability to address reasoning tasks, commonly referred to as chain-of-thought (CoT) (Wei et al., 2022). There have been significant advancements in improving the reasoning abilities of LLMs in terms of end-task performance, such as intricate math word reasoning (Shum et al., 2023; Diao et al., 2023; Zheng et al., 2023; Zhang et al., 2023b), decision-making tasks (Yao et al., 2022), and tasks that involve extensive search and tactical planning (Yao et al., 2023; Besta et al., 2023; Ding et al., 2023). Another line of research lies in the analysis of the reasoning steps themselves with or without human-annotated ones (Clinciu et al., 2021; Welleck et al., 2022; Saparov and He, 2022; Golovneva et al., 2022; Prasad et al., 2023; He et al., 2023).
@@
Wei et al. (2022); Ye and Durrett (2022) have manually inspected whether reasoning steps aligns with the correctness of ultimate answers in both correct and incorrect answer groups through case studies. Moreover, Prasad et al. (2023); He et al. (2023) find that elevating the quality of the rationales could potentially enhance task performance. Since aggregate task performance is a summation of the accuracy of final answers across individual instances, these observations offer an initial qualitative scrutiny of the rationales and their impact on the accuracy of the final predictions. Despite these insights, there remains a shortage of quantitative proof to substantiate whether the evaluation of reasoning chains can reliably affirm the validity of the final outcomes. To bridge this gap, we introduce the R2PE (Relationship of Rationales to Performance Evaluation) benchmark, a test bed designed to quantitatively investigate this question.
@@
The susceptibility of LLMs to generate incorrect information has been underscored by OpenAI-Blog (2022); Zhao et al. (2023a). Meanwhile, Ye and Durrett (2022) associate false predictions with nonfactual explanations. Furthermore, Wei et al. (2022) have demonstrated that, even in arithmetic reasoning tasks, incorrect rationales can occasionally yield correct outcomes, and a variety of errors may occur at intermediate reasoning steps. To perform a qualitative evaluation of rationales with the aim of validating final predictions, the R2PE benchmark is established to integrate a diverse spectrum of reasoning tasks, covering mathematical, commonsense, physical, and textual reasoning (including fact verification and open-domain question answering). These tasks extend across domains such as mathematics, common knowledge, physics, literature, and general world knowledge. Responses are collected from six distinct LLMs to promote a broad generality of the findings. The characteristics and the creation steps of the R2PE benchmark are illustrated in Figure 2.
@@
Each evaluated question or claim begins with the generation of multiple rationales, leading to an answer aggregation based on majority voting (Wang et al., 2022) after extraction and manual cleaning. Once the final outcomes are labeled as true or false using an exact match, our goal is to derive a discernibility score (DS), which is intended as an indicator of the credibility of the final answer. To fully exploit the information contained in all reasoning chains, we present a process discernibility score (PDS) that substantially exceeds the answer discernibility score (ADS) baseline that counts the number of the same answers. As shown in Figure 1, PDS can detect conflicting information about the actor and predict the incorrectness of the final answer. Figure 2 presents a succinct overview of both our benchmark and the associated methodology.
@@
  • We introduce a process discernibility score (PDS) framework that aggregates the information in different chains of reasoning to validate the correctness of the final output.
  • Comprehensive experiments of PDS in R2PE reveal its superiority over the answer discernibility score (ADS), leading to consistent increases in the F1 score and AUC-PR.
  • Further experiments show the effectiveness of our approach in improving open domain question answering performance when combined with verify-and-edit (Zhao et al., 2023a).
@@ 2 R2PE Benchmark
R2PE serves as a comprehensive platform for assessing the reliability of LLM-generated answers in CoT reasoning. R2PE is meticulously constructed with several critical characteristics in mind: (i) Assessability: every instance within R2PE can be verifiable as true or false based on a certain criterion. (ii) Task Diversity: The benchmark should encompass a wide range of reasoning datasets, featuring various answer formats, and spanning different task categories across multiple domains. (iii) Generalizability across LLMs: Responses should be elicited using different LLMs to ensure broad applicability. (iv) High Quality: To minimize instances where correct answers are inaccurately labeled as false due to extraction failures, it is crucial that answers are precisely extracted from responses.
@@ 2.1 Construction Process
Data Collection. we utilize a total of eight datasets including benchmarks for mathematical reasoning such as GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021), common sense reasoning tasks like StrategyQA (Geva et al., 2021) and Play dialogue from BIG-bench collections (Srivastava et al., 2022), physical reasoning tasks (Physics from BIG-bench collections (Srivastava et al., 2022)); fact verification (FEVER (Thorne et al., 2018)) and open-domain question answering (HotpotQA (Yang et al., 2018), 2WikiMultihop (Ho et al., 2020)). Each dataset selected meets the criterion of assessability through verifiable answers. These tasks not only vary in answer formats, including numerical, yes/no, multiple choice, and free form, but also span an array of domains such as mathematics, commonsense, literature, physics, and general world knowledge (see Appendix A.1).
@@
Response Generation. For each question (or claim in the FEVER dataset) $Q$, the LLM is prompted using CoT to produce $n$ responses $(R_{1},R_{2},...,R_{n})$, respectively. The full prompts are given in Appendix F. The responses are aggregated from six distinct LLMs to ensure the generalizability: text-davinci-003 (Ouyang et al., 2022), GPT-3.5-turbo (OpenAI-Blog, 2022), its instruct-trained variant GPT-3.5-turbo-instruct, Gemini Pro, Mixtral-8x7b (Jiang et al., 2024), and mistral-medium. A concise exposition along with detailed settings, is deferred to Appendix A.2.
@@
Extraction. Rationale $r_{i}$ and the corresponding anwser $a_{i}$ need to be isolated from each response $R_{i}$. During the pilot extraction period, we use the prompt cue "The answer is" as a delimiter to segregate responses. Our observations indicated that almost all LLMs yielded some outputs that did not conform to the expected answer format. Given the varying responses styles across different LLMs, the answer trigger words are identified for different LLMs and datasets after observing the original responses to facilitate extraction (see Table 7). This strategy effectively segments most of the responses.
@@
Manual Cleaning. To maintain a high-quality benchmark, manual inspection and cleaning are performed to handle unusual cases. For atypical responses that deviate from the recognized patterns, we either manually separate the response to derive the answer or assign a special marker when separation is not feasible. Additionally, special attention has been paid to responses that are not sentences. A detailed description of the extraction and manual cleaning procedure is provided in Appendix A.3.
@@
Labeling. Upon completion of extraction and manual cleaning, $n$ responses with reasoning chains and corresponding responses $R=\{R_{1}=(r_{1},a_{1}),R_{2}=(r_{2},a_{2}),...,R_{n}=(r_{n},a_{n})\}$ are collected for each question $Q$. With these pairs of rationale answers extracted, we can aggregate the outputs to obtain the final answer by the majority votes: $a=argmax_{a_{i}}freq(a_{i})$, where $freq(a_{i})$ denotes the frequency in which $a_{i}$ appears. The final result $a$ of each question $Q$ is then cross-referenced with its ground truth; A match results in a "TRUE" label, while a mismatch is designated as "FALSE".
@@ 2.2 R2PE Overview
Our R2PE benchmark comprises a diverse collection of data derived from six LLMs on eight reasoning tasks. These tasks encompass a variety of types, answer formats, and domains, making the benchmark rich and comprehensive. Overall, it includes approximately 38.3K instances with around 20.2K labeled as TRUE and 18.1K as FALSE. The structure of R2PE allows for the organization of instances into 45 distinct subsets based on the original dataset and the LLM used to generate responses. Hereinafter, subsets are denoted as (<dataset name> from <LLM name>), with detailed statistics and concrete examples available in Appendix B.
Each instance $e$ within R2PE incorporates a series of structured data fields: question or claim $Q$, the associated dataset name, the queried LLM name, five responses alongside their reasoning chains and answers $R=\{R_{1}=(r_{1},a_{1}),R_{2}=(r_{2},a_{2}),...,R_{5}=(r_{5},a_{5})\}$, the final output $a$, the ground-truth answer of question $Q$, and the label $L\in\{T,F\}$ to indicate whether the generated answer $a$ matches the ground truth. Detailed explications of the constituent data fields and illustrative examples are presented in Appendix B.
With R2PE, we aim to predict the correctness of final outcomes based on the intermediate reasoning steps. The response set $R$ serves as an input for predicting the label $L$ for every instance $e$. We introduce a numerical criterion, the discernibility score (DS), to encapsulate the quality of $R$.A low DS might suggest a potential mismatch between the final output $a$ and the ground-truth answer. Hence, we will classify the example as false: $\hat{L}=F$, if its DS falls below a certain threshold $H$.
@@ 3 Process Discernibility Score
Our goal is to find a suitable DS that can be a good indicator for the veracity of final predictions. We suggest process supervision be taken into account recognizing its potential to overcome the limitations of DS based solely on answer agreement, which tends to overlook a great number of examples with the label $L=F$. To demonstrate its usefulness, we have conducted comprehensive experiments (results are presented in Section 4).
@@ 3.1 The Fallacy of Answer Agreement
Figure 3: Visualization of results on the subset (HotpotQA from GPT-4): ADS (left), PDS (middle), and PDS-ADS (right). ADS has high precision: large answer agreement does not always guarantee accurate predictions. PDS yields desirable outcomes, while PDS-ADS has poor precision.
Based on the insights of Wang et al. (2022); Zhao et al. (2023a), the degree of consensus among answers, is posited as a diagnostic tool to gauge the instances in which LLMs may make wrong predictions. We refer to the answer agreement number as Answer Discernibility Score (ADS) hereinafter. The ADS can thus be adopted as a natural baseline in our R2PE benchmark. A high ADS often suggests a high likelihood that the proposed answer is correct. Given this premise, designating a threshold of $H=n/2$ is a reasonable strategy. If the ADS is below that midpoint, $\hat{L}=F$ is predicted.
However, LLMs may generate substantially consistent but incorrect answers in reasoning tasks. Figure 1 showcases one such example, where the LLM is queried about the nationality of an actor from "Hostage for a Day". The LLM gives the answer "American" four times and "Canadian" once. Here, the ADS stands at 4, which erroneously points to a seemingly accurate but ultimately false consensus. This example serves to warn against overreliance on ADS as the sole indicator of the veracity of the final answer in reasoning tasks.
@@ 3.2 Process Discernibility Score
In our R2PE benchmark, each response $R_{i}$ constitutes a rationale $r_{i}$ paired with an answer $a_{i}$. The ADS, however, limits its analysis to the answers $(a_{1},...,a_{n})$, which takes up only a small proportion of the information contained in $R$. Considering the demonstrated significance of intermediate reasoning steps (Yoran et al., 2023), a DS that makes use of rationales would be preferable.
Inspired by the hypothesis that LLMs tend to generate inaccurate information when they are uncertain about their predictions (Manakul et al., 2023), we introduce PDS framework, which is designed to assess discrepancies among all responses, going beyond explicit verification of answers alone.
To quantify the similarity between two responses, we employ Pairwise Process Supervision Score PPSS(A, B) to measure the degree to which the content of B is both included in and opposes that of A. PPSS is confined within the interval (-1, 1), where a positive score indicates a higher level of similarity relative to contradiction. For a set of n responses, we can calculate the average of PPSS across all ordered response pairs, culminating in the Process Supervision Score (PSS): $\text{PSS}=\text{mean}_{i\neq j}\text{PPSS}(R_{i},R_{j})$. This allows PSS to evaluate the information across all responses. To underscore the final answers, we get
@@
Subsequently, in a manner analogous to the ADS prediction procedure, we classify instances where the PDS falls below the threshold $H$ as $\hat{L}=F$ (indicating a likely falsehood), whereas all remaining instances are designated as $\hat{L}=T$.
Compared to ADS, our PDS is capable of overseeing reasoning processes as well as doing implicit answer checking ($\overline{\text{ADS}}$ term in Equation (1)). As depicted in Figure 1, the PDS can detect contradictive information about the name of the actor playing a Russian hostage taker in "Hostage for a Day” among multiple reasoning chains (highlighted by red) and give a correct label $\hat{L}=F$, while the ADS only assesses the agreements among the answers and leads to a wrong prediction $\hat{L}=T$.

### v4

# mO shizhe[#1]

@@ Abstract
Chain-of-Thought (CoT) prompting has marked a significant advancement in enhancing the reasoning capabilities of large language models (LLMs). Previous studies have developed various extensions of CoT, which focus primarily on enhancing end-task performance. In addition, there has been research on assessing the quality of reasoning chains in CoT. This raises an intriguing question: Is it possible to predict the accuracy of LLM outputs by scrutinizing the reasoning chains they generate? To answer this research question, we introduce a benchmark, R2PE, designed specifically to explore the relationship between reasoning chains and performance in various reasoning tasks spanning five different domains. This benchmark aims to measure the falsehood of the final output of LLMs based on the reasoning steps. To make full use of information in multiple reasoning chains, we propose the process discernibility score (PDS) framework that beats the answer-checking baseline by a large margin. Concretely, this resulted in an average of $5.1\%$ increase in the F1 score and $2.97\%$ improvement in AUC-PR across all 45 subsets within R2PE. We further demonstrate our PDS’s efficacy in advancing open-domain QA accuracy. Codes and data are available at https://github.com/XinXU-USTC/R2PE.git.
Figure 1: An example from HotpotQA that GPT-4 outputs highly consistent but wrong answers. In this particular example, our PDS can detect conflicting information about the actor (colored by red) and predicts that the answer will be incorrect because it is less than zero, while ADS predicts the answer to be correct because it is greater than 2.5.
@@ 1 Introduction
Recent development in large language models (LLMs) (OpenAI, 2023; Bubeck et al., 2023; Yang et al., 2024a) has showcased their remarkable aptitude for tackling diverse downstream tasks. Given several demonstrations with reasoning steps, LLMs exhibit a formidable capability to address reasoning tasks, commonly referred to as chain-of-thought (CoT) (Wei et al., 2022). There have been significant advancements in improving the reasoning abilities of LLMs in terms of end-task performance, such as intricate math word reasoning (Shum et al., 2023; Diao et al., 2023; Zheng et al., 2023; Yu et al., 2023; Xu et al., 2024; Tong et al., 2024; Yang et al., 2024b; Yan et al., 2024), and tasks that involve extensive search and tactical planning (Yao et al., 2023). Another line of research lies in the analysis of the reasoning steps themselves with or without human-annotated ones (Clinciu et al., 2021; Prasad et al., 2023; Xia et al., 2024).
@@
Wei et al. (2022); Ye and Durrett (2022) have manually inspected whether reasoning steps align with the correctness of ultimate answers in both correct and incorrect answer groups through case studies. Moreover, Prasad et al. (2023); He et al. (2023) find that elevating the quality of the rationales could potentially enhance task performance. Since aggregate task performance is a summation of the accuracy of final answers across individual instances, these observations offer an initial qualitative scrutiny of the reasoning chains and their impact on the accuracy of the final predictions. Despite these insights, there remains a shortage of quantitative proof to substantiate whether the evaluation of reasoning chains can reliably affirm the validity of the final outcomes. To bridge this gap, we introduce the R2PE (Relationship of Rationales to Performance Evaluation) benchmark, a test bed designed to quantitatively investigate this question.
@@
The susceptibility of LLMs to generate incorrect information has been underscored by OpenAI-Blog (2022); Zhao et al. (2023). Meanwhile, Ye and Durrett (2022) associate false predictions with nonfactual explanations. Furthermore, Wei et al. (2022) have demonstrated that, even in arithmetic reasoning tasks, incorrect rationales can occasionally yield correct outcomes, and a variety of errors may occur at intermediate reasoning steps. To perform a qualitative evaluation of rationales to validate final predictions, the R2PE benchmark is established to integrate a diverse spectrum of reasoning tasks, covering mathematical, commonsense, physical, and textual reasoning (including fact verification and open-domain question answering). These tasks extend across domains such as mathematics, common knowledge, physics, literature, and general world knowledge. Responses are collected from six distinct LLMs to promote a broad generality of the findings. The characteristics and the creation steps of the R2PE benchmark are illustrated in Figure 2.
@@
Each evaluated question or claim begins with the generation of multiple reasoning chains, leading to an answer aggregation based on majority voting (Wang et al., 2022) after extraction and manual cleaning. Once the final outcomes are labeled as true or false using an exact match, our goal is to derive a discernibility score (DS), which is intended as an indicator of the credibility of the final answer. To fully exploit the information contained in all reasoning chains, we present a process discernibility score (PDS) that substantially exceeds the answer discernibility score (ADS) baseline that counts the number of the same answers. As shown in Figure 1, PDS can detect conflicting information about the actor and predict the incorrectness of the final answer. Figure 2 presents a succinct overview of both our benchmark and our method.
@@
  • We introduce a process discernibility score (PDS) framework that aggregates the information in different reasoning chains for CoT verification.
  • Comprehensive experiments of PDS reveal its superiority over the answer discernibility score (ADS) in predicting final answer correctness, leading to consistent increases in the F1 score and AUC-PR.
  • Further experiments show the effectiveness of our approach in improving open domain question answering performance when combined with verify-and-edit (Zhao et al., 2023).
@@ 2 R2PE Benchmark
R2PE serves as a comprehensive platform for verifying LLM-generated answers in CoT reasoning, which is meticulously constructed with the following critical characteristics: (i) Assessability: every instance within R2PE can be verifiable as true or false based on a certain criterion. (ii) Task Diversity: The benchmark should encompass a wide range of reasoning datasets, featuring various answer formats, and spanning different task categories across multiple domains. (iii) Generalizability across LLMs: Responses should be elicited using different LLMs to ensure broad applicability. (iv) High Quality: To minimize instances where correct answers are inaccurately labeled as false due to extraction failures, it is crucial that answers are precisely extracted from responses.
@@ 2.1 Construction Process
Data Collection. we utilize a total of eight datasets including benchmarks for mathematical reasoning such as GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021), common sense reasoning tasks like StrategyQA (Geva et al., 2021) and Play dialogue from BIG-bench collections (Srivastava et al., 2022), physical reasoning tasks (Physics from BIG-bench collections (Srivastava et al., 2022)); fact verification (FEVER (Thorne et al., 2018)) and open-domain question answering (HotpotQA (Yang et al., 2018), 2WikiMultihop (Ho et al., 2020)). Each selected dataset meets the criterion of assessability through verifiable answers. These tasks not only vary in answer formats, including numerical, yes/no, multiple choice, and free form, but also span an array of domains such as mathematics, commonsense, literature, physics, and general world knowledge (see Appendix A.1).
@@
Response Generation. For each question (or claim in the FEVER dataset) $Q$, the LLM is prompted using CoT to produce $n$ responses $(R_{1},R_{2},...,R_{n})$, respectively. The full prompts are given in Appendix G. The responses are aggregated from six distinct LLMs to ensure the generalizability: text-davinci-003, GPT-3.5-turbo (OpenAI-Blog, 2022), its instruct-trained variant GPT-3.5-turbo-instruct, Gemini Pro, Mixtral-8x7b, and mistral-medium. A concise exposition along with detailed settings, is deferred to Appendix A.2.
@@
Extraction. Rationale $r_{i}$ and the corresponding anwser $a_{i}$ need to be isolated from each response $R_{i}$. During the pilot extraction period, we use the prompt cue ”The answer is” as a delimiter to segregate responses. We find that almost all LLMs yielded some outputs that did not conform to the expected answer format. Given the varying responses styles across different LLMs, the answer trigger words are identified for different LLMs and datasets after observing the original responses to facilitate extraction (see Table 9). This strategy effectively segments most of the responses.
@@
Manual Cleaning. To maintain a high-quality benchmark, manual inspection and cleaning are performed to handle unusual cases. For atypical responses that deviate from the recognized patterns, we either manually separate the response to derive the answer or assign a special marker when separation is not feasible. A detailed description of the extraction and manual cleaning procedure is provided in Appendix A.3.
@@
Labeling. Upon completion of extraction and manual cleaning,$R=\{R_{1}=(r_{1},a_{1}),...,R_{n}=(r_{n},a_{n})\}$ are collected for each question $Q$. Then we can aggregate the outputs to obtain the final answer by the majority votes: $a=argmax_{a_{i}}freq(a_{i})$, where $freq(a_{i})$ denotes the frequency in which $a_{i}$ appears. The final result $a$ of each question $Q$ is then compared with its ground truth; A match results in a ”TRUE” label, while a mismatch is assigned as ”FALSE”.
@@ 2.2 R2PE Overview
Our R2PE benchmark comprises a diverse collection of data derived from six LLMs on eight reasoning tasks. These tasks encompass a variety of types, answer formats, and domains, making the benchmark rich and comprehensive. In total, it includes approximately 38.3K instances with around 20.2K labeled as TRUE and 18.1K as FALSE. The structure of R2PE allows for the organization of instances into 45 distinct subsets based on the original dataset and the LLM used to generate responses. Hereinafter, subsets are denoted as (dataset name from LLM name), with detailed statistics and concrete examples available in Appendix B.
Each instance $e$ within R2PE incorporates a series of structured data fields: question or claim $Q$, the associated dataset name, the queried LLM name, five responses alongside their rationales and answers $R=\{R_{1}=(r_{1},a_{1}),R_{2}=(r_{2},a_{2}),...,R_{5}=(r_{5},a_{5})\}$, the final output $a$, the ground-truth answer of question $Q$, and the label $L\in\{T,F\}$ to indicate whether the generated answer $a$ matches the ground truth. Detailed explications of the constituent data fields and illustrative examples are presented in Appendix B.
We aim to predict the correctness of final outcomes based on the intermediate reasoning steps. The response set $R$ serves as an input for predicting the label $L$ for every instance $e$. We introduce a numerical criterion, the discernibility score (DS), to encapsulate the quality of $R$.A low DS might suggest a potential mismatch between the final output $a$ and the ground-truth answer. Hence, we will classify the example as false: $\hat{L}=F$, if its DS falls below a certain threshold $H$.
@@ 3.1 The Fallacy of Answer Agreement
Figure 3: Results on the subset (HotpotQA from GPT-4): ADS (left), PDS (middle), and PDS-ADS (right). ADS has high precision: large answer agreement does not always guarantee accurate predictions. PDS yields desirable outcomes, while PDS-ADS has poor precision.
From Wang et al. (2022); Zhao et al. (2023), the degree of consensus among answers, is posited as a diagnostic tool to gauge the instances in which LLMs may make wrong predictions. We refer to the answer agreement number as Answer Discernibility Score (ADS) hereinafter. A high ADS often suggests a high likelihood that the proposed answer is correct. If the ADS is below that midpoint (i.e., the threshold $H=n/2$), $\hat{L}=F$ is predicted.
However, LLMs may generate substantially consistent but incorrect answers in reasoning tasks. Figure 1 showcases one such example, where the LLM is queried about the nationality of an actor from ”Hostage for a Day”. The LLM gives the answer ”American” four times and ”Canadian” once. Here, the ADS stands at 4, which erroneously points to a seemingly accurate but ultimately false consensus. This example serves to warn against overreliance on ADS as the sole indicator of the veracity of the final answer in reasoning tasks.
@@ 3.2 Process Discernibility Score
In our R2PE benchmark, each response $R_{i}$ constitutes a rationale $r_{i}$ paired with an answer $a_{i}$. The ADS, however, limits its analysis to the answers $(a_{1},...,a_{n})$, which takes up only a small proportion of the information contained in $R$. Inspired by (Yoran et al., 2023), a DS that makes use of rationales would be preferable.
To quantify the similarity between two responses, we employ Pairwise Process Supervision Score PPSS($R_{k}$, $R_{l}$) to measure the degree to which the content of $R_{l}$ is both included in and opposes that of $R_{k}$. PPSS is confined within the interval (-1, 1), where a positive score indicates a higher level of similarity relative to contradiction. For a set of n responses, we can calculate the average of PPSS across all ordered response pairs, culminating in the Process Supervision Score (PSS): $\text{PSS}=\text{mean}_{i\neq j}\text{PPSS}(R_{i},R_{j})$. This allows PSS to evaluate the information across all responses. To underscore the final answers, we get
@@
Subsequently, we classify instances where the PDS falls below the threshold $H$ as $\hat{L}=F$ , whereas all remaining instances are designated as $\hat{L}=T$.
Compared to ADS, our PDS is capable of overseeing reasoning processes as well as doing implicit answer checking ($\overline{\text{ADS}}$ term in Equation (1)). As depicted in Figure 1, the PDS can detect contradictive information about the name of the actor playing a Russian hostage taker in ”Hostage for a Day” among multiple reasoning chains (highlighted by red) and give a correct label $\hat{L}=F$, while the ADS only assesses the agreements among the answers.